Logistics Solutions provides order fulfillment services for dot.com merchants. The company maintains warehouses that stock items carried by its dot.com clients. When a client receives an order from a customer, the order is forwarded to Logistics Solutions, which pulls the item from storage, packs it, and ships it to the customer. The company uses a predetermined variable overhead rate based on direct labor-hours. In the most recent month, 110,000 items were shipped to customers using 3,500 direct labor-hours. The company incurred a total of $9,450 in variable overhead costs. According to the company’s standards, 0.03 direct labor-hours are required to fulfill an order for one item and the variable overhead rate is $2.75 per direct labor-hour. Required: 1. What is the standard labor-hours allowed (SH) to ship 110,000 items to customers? 2. What is the standard variable overhead cost allowed (SH × SR) to ship 110,000 items to customers? 3. What is the variable overhead spending variance? 4. What is the variable overhead rate variance and the variable overhead efficiency variance?

1.Standard quantity of labor-hours allowed2.Standard variable overhead cost allowed3.Variable overhead spending variance4.Variable overhead rate varianceVariable overhead efficiency variance
Answer #1

Standard hours allowed = 110,000*0.03

= 3300

Standard cost allowed = 3300*2.75

= 9075

Voh spending variance = 9075 - 9450

= 375 U

Voh rate variance = (SR-AR) *AH = (2.75 - 9450/3500),3500

= 175 F

Voh efficiency variance = (SH-AH) *SR = (3300-3500)*2.75

= 550 U